# An entropy formula for the Deep Linear Network

### v1

@@ Abstract.
We study the Riemannian geometry of the Deep Linear Network (DLN) as a foundation for a thermodynamic description of the learning process. The main tools are the use of group actions to analyze overparametrization and the use of Riemannian submersion from the space of parameters to the space of observables. The foliation of the balanced manifold in the parameter space by group orbits is used to define and compute a Boltzmann entropy. We also show that the Riemannian geometry on the space of observables defined in [2] is obtained by Riemannian submersion of the balanced manifold. The main technical step is an explicit construction of an orthonormal basis for the tangent space of the balanced manifold using the theory of Jacobi matrices.
@@ 1. Overview
The Deep Linear Network (DLN) is a phenomenological model of deep learning based on the composition of linear functions. It was introduced by Arora, Cohen and Hazan in 2017 [1]. The DLN builds on earlier models in neural computing of a similar nature (in particular [3]), but differs in its emphasis on the role of depth.
In this paper, and the companion papers [24, 27], we use the DLN to develop a rigorous thermodynamic description of the learning process. The main novelty in our work is the systematic use of the geometric theory of dynamical systems. This is in contrast with the more common approach based on statistical learning theory. We study both the Riemannian and symplectic geometry that underlie training dynamics and use it as the foundation for a thermodynamic description.
This paper develops the Riemannian geometry. We provide explicit descriptions of the underlying Riemannian metrics and define a natural Boltzmann entropy. The related stochastic dynamics are described by a Riemannian Langevin equation (RLE) which is presented in the companion paper [27]. Both these papers rely on analogies with random matrix theory. The determinantal formula for the entropy in this paper (Theorem 4) is a variant of well-known formulas in random matrix theory. The RLE for the DLN in [27] is a geometric stochastic flow analogous to Dyson Brownian motion. In the companion paper [24], we develop the symplectic geometry of the DLN and provide a variational characterization of balancedness. This result is obtained by developing an analogy with Kalman’s minimal realizability theory in linear systems theory. Both these analogies reveal unexpected mathematical structure in the DLN and new directions for study.
@@
The use of group actions and Riemannanian geometry builds on earlier attempts in control and systems theory, especially by Brockett and his co-authors, to develop gradient flows that solve combinatorial optimization problems and logical tasks. Brockett’s work was also motivated by the study of neural networks and blurs the usual distinction between analog and digital computing [7, 8, 9]. The reappearance of Kalman’s realizability theory in the DLN reveals the continued utility of geometric control theory in the analysis of neural networks.
The Riemannian geometry studied here is of independent mathematical interest. An explicit description of geodesics in the DLN geometry, based on the classical theory of confocal quadrics and Theorem 8 below, is presented by Chen [10]. His work generalizes Brenier’s formula for geodesics in the Bures-Wasserstein metric to the DLN, providing a new connection between optimal transport theory and deep learning. The entropy formula in this paper has also been used to study equilibrium measures in the DLN [11], formalizing the notion of entropic regularization and selection of minimizers by small noise in deep learning. A surprising feature of the DLN, in contrast with random matrix theory, is that the equilibrium measures do not show repulsion between singular values. Finally, the use of Riemannian gradient flows also provides a certain unity between interior point methods for convex optimization and deep learning as discussed in [26].
@@ 2. Background
We begin by reviewing rigorous results by other authors that have stimulated our work. The expository article [25] provides more details on sections 2.22.4 below.
@@ 2.4. Previous results
Theorem 1 (Arora, Cohen, Hazan [1]).
@@
Theorem 2 (Arora, Cohen, Hazan [1]).
@@
Theorem 3 (Bah, Rauhut, Terstiege, Westdickenberg [2]).
@@ 3.2. Fibers, group orbits and the Boltzmann entropy
where the volume form is with respect to the metric inherited by $\mathcal{M}$ because of its embedding in $\mathbb{M}_{d}^{N}$. We compute this volume form, and thus $S(X)$, exactly using methods from random matrix theory (see Theorem 4). In the companion paper [27] we introduce a Riemannian Langevin equation (RLE) for the DLN, which is a natural stochastic perturbation of the gradient flow (2.6). This construction is based on an analogy with Dyson Brownian motion and provides microscopic stochastic dynamics that are consistent with the entropy formula below.
@@ 3.3. The entropy formula
When $X$ has full rank but repeated singular values, $S(X)$ may be evaluated by taking limits in the above formula using L’Hospital’s rule. The value of the constant $c_{d}$ is $2^{\frac{1}{2}d(d+3)}\prod_{r=1}^{d}\frac{\pi^{\frac{r}{2}}}{\Gamma(\frac{r}{2})}$ [29].
@@
In the companion paper [24], we show that $\mathbf{G}$ are obtained from a moment map on $\mathcal{F}_{X}$ associated to the action of the subgroup $O_{d}^{N-1}\subset GL(d)^{N-1}$. This approach yields the general setting of the entropy, with
@@ 3.4. Riemannian submersion
The metric $g^{N}$ was first introduced in [2] using the linear operator $\mathcal{A}_{N,X}$ defined in equations (2.19)–(2.20). On the other hand, the balanced manifold $\mathcal{M}$ inherits a Riemannian metric from its embedding in $\mathbb{M}_{d}^{N}$; we denote the resulting Riemannian manifold by $(\mathcal{M},\iota)$. We prove
@@
Theorem 8 provides the natural geometric explanation for certain unexpected properties of the metric $g^{N}$ observed in the literature. Let us motivate Theorem 8 by recalling two such properties: the diagonalization of $\mathcal{A}_{N,X}$ using the SVD and a determinantal formula related to Theorem 4 [15].
@@
The proof is a direct calculation (see [25, §7] for details). Every matrix $Z\in T_{X}\mathfrak{M}_{d}$ may be expressed in this basis as a sum $Z=\sum_{k,l=1}^{d}Z_{kl}q_{N,k}q_{0,l}^{T}$. Equation (2.20) and Lemma 1 then imply
@@
Theorem 9 (Cohen, Menon, Veraszto [15]).
@@ 4.1. Parametrizing the balanced variety $\mathcal{M}_{0}$
When $X$ has full rank and the singular values are distinct, then each $W_{p}$ also has full rank and distinct singular values. Now the SVD of each $W_{p}$ is unique except upto a permutation and the singular vectors depend analytically on $W_{p}$ by the standard perturbation theory for eigenvalues [21]. Thus, $\mathfrak{z}$ is locally an analytic diffeomorphism. ∎
@@ 5.3. Proofs of Lemmas
The formulas for the eigenvalues and eigenvectors of $C$ are well known and are easily modified to yield the identities (5.9)–(5.12) in Lemma 9 (see for example [17, p.476]). The reader unfamiliar with this theory may simply verify these identities to complete the proof. ∎
@@ 7.2. Implicit bias
In deep learning, the function $f$ is constructed using neural networks whose parameters are obtained by using gradient descent to minimize the empirical loss function constructed from the data. The fundamental mystery in deep learning is why deep neural networks generalize so well, without overfitting, even in the absence of an explicit regularizer. A striking dynamic demonstration of the distinction between the predictions of (traditional) statistical learning theory and deep learning is provided by the double-descent curve of Belkin et al [4, 5]. This ability of deep neural networks to generalize without overfitting is known as implicit bias or implicit regularization. We refer the reader to a recent review by Vardi that summarizes several approaches to implicit bias [30].
The main conjecture that underlies our work is that implicit bias in deep learning is a dynamic phenomena that originates in the geometry of gradient fields for overparametrized functions. To this end, we have focused on a rigorous analysis of the geometric consequences of overparametrization in the DLN using Lie groups. We have identified robust properties of the gradient flow that do not rely on the form of the loss function or on simplifying assumptions on the dimension and depth. The main advantages of our approach are that it provides a deeper understanding of the fundamental concept of balancedness, it reveals the intrinsic Riemannian geometry of the learning process, it allows the inclusion of noise as a selection principle, and that it provides exact formulas (such as the entropy formula) without any restrictions on $d$ and $N$. We note that there have been other recent attempts to derive effective theories for deep learning using geometric methods (see in particular [12]).
The mathematical foundation of our approach is dynamical systems theory, not statistical learning theory. But there are, of course, several places where the methods overlap. In addition to the prior work [1, 2] whose importance we have discussed in Section 3, there have been several attempts to understand implicit bias in linear networks. These include a characterization under simplifying assumptions (with margin-based generalization bounds for logistic regression when $d=1$ [20] and with the nuclear norm when $N=2$ [18, 19]). An interesting recent modification of linear networks are the spectral neural networks introduced in [14]. There have also been several studies of the convergence to critical points in the DLN for different choices of learning tasks [6, 16, 28] or simplifying assumptions on dimension (e.g. $N=3$ and $d=\infty$ in [13]). There has also been rigorous geometric analysis of the expressivity properties of linear convolutional networks [22, 23]. It is of interest to combine this geometric analysis with gradient dynamics.
@@ 7.3. Entropic regularization
The entropy formula may be used to regularize the gradient flow by modifying the gradient dynamics to include small noise in the gauge group $O_{d}^{N-1}$. While the rigorous theory is developed in [27], we describe the main insights here.
@@
We show in [27] (see also [25, §12] for a simplified exposition) that the gradient dynamics of $X$ correspond to stochastic dynamics for $\mathbf{W}$ on the balanced manifold $\mathcal{M}$ described by a Riemannian Langevin equation (RLE). The stochastic forcing of this RLE is Brownian motion on $\mathcal{O}_{X}$ and an explicit description of this Brownian motion is provided by the orthonormal basis computed in Theorem 11.
It is shown in [11] that the entropy is concave (in the Euclidean geometry on $\mathbb{M}_{d})$. Thus, the inclusion of the entropy in the free energy regularizes the loss function $E(X)$ allowing us to characterize the minimizers of $F_{\beta}$ for certain loss functions $E$. While the entropy formula was discovered through an analogy with random matrix theory (RMT), a surprising feature is that the minimizers of the free energy for the DLN do not show singular value repulsion.
@@
Partial results in this direction are presented in [11].
@@ 7.4. Equilibrium thermodynamics in the DLN
Thus, the important mathematical structure that allows us to treat the physical and information theoretic entropy in the same manner is Brownian motion on a manifold. In the setting of the DLN, Brownian motion in $(\mathcal{O}_{X},\iota)$ provides a precise description of fluctuations in the gauge group $O_{d}^{N-1}$. This ensemble of paths is our thermodynamic system and it is conveniently visualized through Riemannian submersion. As $X$ evolves downstairs according to the gradient flow (7.2), the ensemble of paths upstairs evolves in a quasistatic manner. This physical caricature, along with its rigorous mathematical formulation, is described in detail in the forthcoming paper [27].

### v2

@@ Abstract.
We study the Riemannian geometry of the Deep Linear Network (DLN) as a foundation for a thermodynamic description of the learning process. The main tools are the use of group actions to analyze overparametrization and the use of Riemannian submersion from the space of parameters to the space of observables. The foliation of the balanced manifold in the parameter space by group orbits is used to define and compute a Boltzmann entropy. We also show that the Riemannian geometry on the space of observables defined in [Bah] is obtained by Riemannian submersion of the balanced manifold. The main technical step is an explicit construction of an orthonormal basis for the tangent space of the balanced manifold using the theory of Jacobi matrices.
@@ 1. Overview
The Deep Linear Network (DLN) is a phenomenological model of deep learning based on the composition of linear functions. It was introduced by Arora, Cohen and Hazan in 2017 [ACH]. The DLN builds on earlier models in neural computing of a similar nature (in particular [Baldi-Hornik]), but differs in its emphasis on the role of depth.
In this paper, and the companion papers [LM1, MY-RLE], we use the DLN to develop a rigorous thermodynamic description of the learning process. The main novelty in our work is the systematic use of the geometric theory of dynamical systems. This is in contrast with the more common approach based on statistical learning theory. We study both the Riemannian and symplectic geometry that underlie training dynamics and use it as the foundation for a thermodynamic description.
This paper develops the Riemannian geometry. We provide explicit descriptions of the underlying Riemannian metrics and define a natural Boltzmann entropy. The related stochastic dynamics are described by a Riemannian Langevin equation (RLE) which is presented in the companion paper [MY-RLE]. Both these papers rely on analogies with random matrix theory. The determinantal formula for the entropy in this paper (Theorem 4) is a variant of well-known formulas in random matrix theory. The RLE for the DLN in [MY-RLE] is a geometric stochastic flow analogous to Dyson Brownian motion. In the companion paper [LM1], we develop the symplectic geometry of the DLN and provide a variational characterization of balancedness. This result is obtained by developing an analogy with Kalman’s minimal realizability theory in linear systems theory. Both these analogies reveal unexpected mathematical structure in the DLN and new directions for study.
@@
The use of group actions and Riemannanian geometry builds on earlier attempts in control and systems theory, especially by Brockett and his co-authors, to develop gradient flows that solve combinatorial optimization problems and logical tasks. Brockett’s work was also motivated by the study of neural networks and blurs the usual distinction between analog and digital computing [Brockett-transient, Brockett, Brockett-automata]. The reappearance of Kalman’s realizability theory in the DLN reveals the continued utility of geometric control theory in the analysis of neural networks.
The Riemannian geometry studied here is of independent mathematical interest. An explicit description of geodesics in the DLN geometry, based on the classical theory of confocal quadrics and Theorem 8 below, is presented by Chen [Chen1]. His work generalizes Brenier’s formula for geodesics in the Bures-Wasserstein metric to the DLN, providing a new connection between optimal transport theory and deep learning. The entropy formula in this paper has also been used to study equilibrium measures in the DLN [Chen2], formalizing the notion of entropic regularization and selection of minimizers by small noise in deep learning. A surprising feature of the DLN, in contrast with random matrix theory, is that the equilibrium measures do not show repulsion between singular values. Finally, the use of Riemannian gradient flows also provides a certain unity between interior point methods for convex optimization and deep learning as discussed in [MY-conic].
@@ 2. Background
We begin by reviewing rigorous results by other authors that have stimulated our work. The expository article [GM-dln] provides more details on sections 2.22.4 below.
@@ 2.4. Previous results
Theorem 1 (Arora, Cohen, Hazan [ACH]).
@@
Theorem 2 (Arora, Cohen, Hazan [ACH]).
@@
Theorem 3 (Bah, Rauhut, Terstiege, Westdickenberg [Bah]).
@@ 3.2. Fibers, group orbits and the Boltzmann entropy
where the volume form is with respect to the metric inherited by $\mathcal{M}$ because of its embedding in $\mathbb{M}_{d}^{N}$. We compute this volume form, and thus $S(X)$, exactly using methods from random matrix theory (see Theorem 4). In the companion paper [MY-RLE] we introduce a Riemannian Langevin equation (RLE) for the DLN, which is a natural stochastic perturbation of the gradient flow (2.6). This construction is based on an analogy with Dyson Brownian motion and provides microscopic stochastic dynamics that are consistent with the entropy formula below.
@@ 3.3. The entropy formula
When $X$ has full rank but repeated singular values, $S(X)$ may be evaluated by taking limits in the above formula using L’Hospital’s rule. The value of the constant $c_{d}$ is $2^{\frac{1}{2}d(d+3)}\prod_{r=1}^{d}\frac{\pi^{\frac{r}{2}}}{\Gamma(\frac{r}{2})}$ [Ponting].
@@
In the companion paper [LM1], we show that $\mathbf{G}$ are obtained from a moment map on $\mathcal{F}_{X}$ associated to the action of the subgroup $O_{d}^{N-1}\subset GL(d)^{N-1}$. This approach yields the general setting of the entropy, with
@@ 3.4. Riemannian submersion
The metric $g^{N}$ was first introduced in [Bah] using the linear operator $\mathcal{A}_{N,X}$ defined in equations (2.19)–(2.20). On the other hand, the balanced manifold $\mathcal{M}$ inherits a Riemannian metric from its embedding in $\mathbb{M}_{d}^{N}$; we denote the resulting Riemannian manifold by $(\mathcal{M},\iota)$. We prove
@@
Theorem 8 provides the natural geometric explanation for certain unexpected properties of the metric $g^{N}$ observed in the literature. Let us motivate Theorem 8 by recalling two such properties: the diagonalization of $\mathcal{A}_{N,X}$ using the SVD and a determinantal formula related to Theorem 4 [CMV].
@@
The proof is a direct calculation (see [GM-dln, §7] for details). Every matrix $Z\in T_{X}\mathfrak{M}_{d}$ may be expressed in this basis as a sum $Z=\sum_{k,l=1}^{d}Z_{kl}q_{N,k}q_{0,l}^{T}$. Equation (2.20) and Lemma 1 then imply
@@
Theorem 9 (Cohen, Menon, Veraszto [CMV]).
@@ 4.1. Parametrizing the balanced variety $\mathcal{M}_{0}$
When $X$ has full rank and the singular values are distinct, then each $W_{p}$ also has full rank and distinct singular values. Now the SVD of each $W_{p}$ is unique except upto a permutation and the singular vectors depend analytically on $W_{p}$ by the standard perturbation theory for eigenvalues [Kato]. Thus, $\mathfrak{z}$ is locally an analytic diffeomorphism. ∎
@@ 5.3. Proofs of Lemmas
The formulas for the eigenvalues and eigenvectors of $C$ are well known and are easily modified to yield the identities (5.9)–(5.12) in Lemma 9 (see for example [GvL, p.476]). The reader unfamiliar with this theory may simply verify these identities to complete the proof. ∎
@@ 7.2. Implicit bias
In deep learning, the function $f$ is constructed using neural networks whose parameters are obtained by using gradient descent to minimize the empirical loss function constructed from the data. The fundamental mystery in deep learning is why deep neural networks generalize so well, without overfitting, even in the absence of an explicit regularizer. A striking dynamic demonstration of the distinction between the predictions of (traditional) statistical learning theory and deep learning is provided by the double-descent curve of Belkin et al [Belkin-dd, Belkin-dd2]. This ability of deep neural networks to generalize without overfitting is known as implicit bias or implicit regularization. We refer the reader to a recent review by Vardi that summarizes several approaches to implicit bias [Vardi].
The main conjecture that underlies our work is that implicit bias in deep learning is a dynamic phenomena that originates in the geometry of gradient fields for overparametrized functions. To this end, we have focused on a rigorous analysis of the geometric consequences of overparametrization in the DLN using Lie groups. We have identified robust properties of the gradient flow that do not rely on the form of the loss function or on simplifying assumptions on the dimension and depth. The main advantages of our approach are that it provides a deeper understanding of the fundamental concept of balancedness, it reveals the intrinsic Riemannian geometry of the learning process, it allows the inclusion of noise as a selection principle, and that it provides exact formulas (such as the entropy formula) without any restrictions on $d$ and $N$. We note that there have been other recent attempts to derive effective theories for deep learning using geometric methods (see in particular [t-chen]).
The mathematical foundation of our approach is dynamical systems theory, not statistical learning theory. But there are, of course, several places where the methods overlap. In addition to the prior work [ACH, Bah] whose importance we have discussed in Section 3, there have been several attempts to understand implicit bias in linear networks. These include a characterization under simplifying assumptions (with margin-based generalization bounds for logistic regression when $d=1$ [Telgarsky] and with the nuclear norm when $N=2$ [Gunasekar2018, Gunasekar2017]). An interesting recent modification of linear networks are the spectral neural networks introduced in [ghosh]. There have also been several studies of the convergence to critical points in the DLN for different choices of learning tasks [Brechet-Montufar, Ge1, Rauhut2] or simplifying assumptions on dimension (e.g. $N=3$ and $d=\infty$ in [Figalli]). There has also been rigorous geometric analysis of the expressivity properties of linear convolutional networks [Montufar1, Montufar2]. It is of interest to combine this geometric analysis with gradient dynamics.
@@ 7.3. Entropic regularization
The entropy formula may be used to regularize the gradient flow by modifying the gradient dynamics to include small noise in the gauge group $O_{d}^{N-1}$. While the rigorous theory is developed in [MY-RLE], we describe the main insights here.
@@
We show in [MY-RLE] (see also [GM-dln, §12] for a simplified exposition) that the gradient dynamics of $X$ correspond to stochastic dynamics for $\mathbf{W}$ on the balanced manifold $\mathcal{M}$ described by a Riemannian Langevin equation (RLE). The stochastic forcing of this RLE is Brownian motion on $\mathcal{O}_{X}$ and an explicit description of this Brownian motion is provided by the orthonormal basis computed in Theorem 11.
It is shown in [Chen2] that the entropy is concave (in the Euclidean geometry on $\mathbb{M}_{d})$. Thus, the inclusion of the entropy in the free energy regularizes the loss function $E(X)$ allowing us to characterize the minimizers of $F_{\beta}$ for certain loss functions $E$. While the entropy formula was discovered through an analogy with random matrix theory (RMT), a surprising feature is that the minimizers of the free energy for the DLN do not show singular value repulsion.
@@
Partial results in this direction are presented in [Chen2].
@@ 7.4. Equilibrium thermodynamics in the DLN
Thus, the important mathematical structure that allows us to treat the physical and information theoretic entropy in the same manner is Brownian motion on a manifold. In the setting of the DLN, Brownian motion in $(\mathcal{O}_{X},\iota)$ provides a precise description of fluctuations in the gauge group $O_{d}^{N-1}$. This ensemble of paths is our thermodynamic system and it is conveniently visualized through Riemannian submersion. As $X$ evolves downstairs according to the gradient flow (7.2), the ensemble of paths upstairs evolves in a quasistatic manner. This physical caricature, along with its rigorous mathematical formulation, is described in detail in the forthcoming paper [MY-RLE].